### v1

# Complexity of codes for Ramsey positive sets

Abstract.
Sabok showed that the set of codes for $G_{\delta}$ Ramsey positive subsets of $[\omega]^{\omega}$ is $\mathbf{\Sigma}^{1}_{2}$-complete. We extend this result by providing sufficient conditions for the set of codes for $G_{\delta}$ Ramsey positive subsets of an arbitrary topological Ramsey space to be $\mathbf{\Sigma}^{1}_{2}$-complete.
1. Introduction
A well-known theorem of Ramsey states that given any $k<\omega$ and any $\mathcal{X}\subseteq[\omega]^{k}$, there is an infinite set $A\subseteq\omega$ such that $[A]^{k}\subseteq\mathcal{X}$ or $[A]^{k}\cap\mathcal{X}=\emptyset$. This result fails for $k=\omega$ when the Axiom of Choice is assumed. However, there is a topological characterization of which sets $\mathcal{X}\subseteq[\omega]^{\omega}$ satisfy a stronger related property. Recall that the Ellentuck topology on $[\omega]^{\omega}$ is generated by all sets of the form
$[s,A]=\{B\in[\omega]^{\omega}\mid s\sqsubseteq B\wedge B\subseteq A\}$
where $s\in[\omega]^{<\omega}$, $A\in[\omega]^{\omega}$, and $s\sqsubseteq B$ means that $s$ is an initial segment of $B$. A set $\mathcal{X}\subseteq[\omega]^{\omega}$ is Ramsey if for all non-empty $[s,A]$, there is $B\in[s,A]$ such that $[s,B]\subseteq\mathcal{X}$ or $[s,B]\cap\mathcal{X}=\emptyset$. A set is Ramsey null if the latter always holds, and Ramsey positive otherwise. Silver showed that every analytic set is Ramsey [Sil70]. In fact, Ellentuck proved a stronger result: a set is Ramsey iff it has the Baire property in the Ellentuck topology [Ell74]. We discuss these definitions and analogous results in a more general setting in Section 3.
Sabok proved that the set of codes for $G_{\delta}$ (i.e., $\mathbf{\Pi}^{0}_{2}$) Ramsey positive sets in $[\omega]^{\omega}$ is $\mathbf{\Sigma}^{1}_{2}$-complete [Sab12, Theorem 1]; see Section 3 for the relevant definitions. In the same paper, Sabok proved a general theorem about $\mathbf{\Sigma}^{1}_{2}$-complete sets that has been used in several recent proofs of $\mathbf{\Sigma}^{1}_{2}$-completeness [Sab12, Theorem 2]. For instance, Todorčević and Vidnyánszky showed that the set of closed subgraphs of the shift graph on $[\omega]^{\omega}$ that have finite Borel chromatic number is $\mathbf{\Sigma}^{1}_{2}$-complete [TV21, Theorem 1.3]. Brandt, Chang, Grebík, Grunau, Rozhoň, and Vidnyánszky showed a similar result for graphs of bounded degree: for any $d>2$, the set of Borel acyclic $d$-regular graphs with Borel chromatic number at most $d$ is $\mathbf{\Sigma}^{1}_{2}$-complete [BCG+21, Theorem 1.2]. Finally, Thornton proved that a more general class of Borel constraint satisfaction problems is $\mathbf{\Sigma}^{1}_{2}$-complete, including several other examples from Borel combinatorics [Tho22, Theorem 1.7].
To prove [Sab12, Theorem 1], Sabok considered the set of codes corresponding to a universal $G_{\delta}$ set formed by viewing every $x\in 2^{\omega}$ as representing a countable sequence of trees. Under certain conditions, we show that we can find a continuous reduction from this set of codes to the analogous set of codes associated with a topological Ramsey space satisfying axioms specified by Todorčević in [Tod10]. In particular,
Theorem 1.1.
Suppose $(\mathcal{R},\leq,r)$ satisfies A.1-A.4 from [Tod10, Section 5.1], $\mathcal{R}$ is closed, $\mathcal{AR}$ is countable, and there is a function $m:\mathcal{AR}\to\omega$ satisfying Property 5.1 (see Section 5). Then the set of codes for $G_{\delta}$ Ramsey positive subsets of $\mathcal{R}$ is $\mathbf{\Sigma}^{1}_{2}$-complete.
It turns out that many topological Ramsey spaces satisfy the conditions of this result. We present a few examples in Section 6.
2. Acknowledgments
Thank you to Clinton Conley, Natasha Dobrinen, and Aristotelis Panagiotopoulos for helpful discussions. The author was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2140739, and by the ARCS Foundation.
3. Background
Throughout, we will let $(\mathcal{R},\leq,r)$ denote a triple where $\mathcal{R}$ is a non-empty set, $\leq$ is a reflexive and transitive relation on $\mathcal{R}$, and $r:\mathcal{R}\times\omega\to\mathcal{AR}$ is a map into a set of finite approximations of $\mathcal{R}$. We assume that $\mathcal{AR}$ is countable. In addition, we assume that $(\mathcal{R},\leq,r)$ satisfies axioms A.1, A.2, A.3, and A.4 from [Tod10, Section 5.1] and that $\mathcal{R}$ is closed as a subspace of $\mathcal{AR}^{\omega}$, where $\mathcal{AR}$ is given the discrete topology. Unless otherwise stated, we use $s,t,\ldots$ to denote elements of $\mathcal{AR}$ and $A,B,\ldots$ to denote elements of $\mathcal{R}$.
We write $r_{n}(A)$ to mean $r(A,n)$ for $A\in\mathcal{R}$ and $n<\omega$. For $s\in\mathcal{AR}$ and $A\in\mathcal{R}$, define $s\sqsubseteq A$ iff $s=r_{n}(A)$ for some $n<\omega$. For $s,t\in\mathcal{AR}$, define $s\sqsubseteq t$ iff there are $A\in\mathcal{R}$ and $m\leq n<\omega$ such that $s=r_{m}(A)$ and $t=r_{n}(A)$.
We make use of two different topologies on $\mathcal{R}$. We call the topology induced as a subspace of $\mathcal{AR}^{\omega}$ the metrizable topology. Unless otherwise specified, all topological notions are taken to be in the metrizable topology. Note that the metrizable topology on $\mathcal{R}$ is Polish since $\mathcal{AR}$ is countable and $\mathcal{R}$ is closed. Let
$[s,A]:=\{B\in\mathcal{R}\mid s\sqsubseteq B\wedge B\leq A\}.$
We call the topology generated by all $[s,A]$ the Ellentuck topology.
A set $\mathcal{X}\subseteq\mathcal{R}$ is Ramsey if for every $[s,A]\neq\emptyset$, there is $B\in[s,A]$ such that $[s,B]\subseteq\mathcal{X}$ or $[s,B]\cap\mathcal{X}=\emptyset$. We say that $\mathcal{X}$ is Ramsey null if for every $[s,A]\neq\emptyset$, there is $B\in[s,A]$ such that $[s,B]\cap\mathcal{X}=\emptyset$. If $\mathcal{X}$ is not Ramsey null, then we say that $\mathcal{X}$ is Ramsey positive. Note that a Ramsey positive set need not be Ramsey in general.
Todorčević proved the following general connection between the Ellentuck topology and these Ramsey-theoretic notions [Tod10, Theorem 5.4]:
Theorem 3.1 (Todorčević).
Under the assumptions above, every subset of $\mathcal{R}$ with the Baire property in the Ellentuck topology is Ramsey, and every Ellentuck meager set is Ramsey null.
Since $\mathcal{R}$ is Polish, every analytic subset of $\mathcal{R}$ is Suslin-measurable, hence Ramsey by [Tod10, Corollary 5.9]. Then for every analytic $\mathcal{X}\subseteq\mathcal{R}$, we have that $\mathcal{X}$ is Ramsey positive iff $\mathcal{X}$ contains some $[s,A]\neq\emptyset$.
Recall that $([\omega]^{\omega},\subseteq,r)$ satisfies the assumptions above, where $r_{n}$ is the function mapping an element of $[\omega]^{\omega}$ to the set of its least $n$ elements. We take $s^{\prime},t^{\prime},\ldots$ to denote elements of $[\omega]^{<\omega}$ and $A^{\prime},B^{\prime},\ldots$ to denote elements of $[\omega]^{\omega}$ unless stated otherwise.
Let $\mathbf{\Gamma}$ be a pointclass, $\mathcal{A}\subseteq\mathbf{\Gamma}(X)$, and $U\subseteq\omega^{\omega}\times X$ a universal $\mathbf{\Gamma}$ set. We call $\{x\in\omega^{\omega}\mid U_{x}\in\mathcal{A}\}$ the set of codes for $\mathcal{A}$ in $X$. We will primarily be interested in the set of codes for $G_{\delta}$ Ramsey positive sets in a topological Ramsey space $\mathcal{R}$ as above.
4. The set of codes for $G_{\delta}$ Ramsey positive sets is $\mathbf{\Sigma}^{1}_{2}$
[…]
$G_{x}=\mathcal{R}\setminus\bigcup_{n<\omega}[X_{n}],$
where $[X_{n}]=\{A\in\mathcal{R}\mid\forall k<\omega~{}r_{k}(A)\in X_{n}\}$. Otherwise, let $G_{x}=\emptyset$. Then every $G_{\delta}$ subset of $\mathcal{R}$ is realized as a section of $G$. Define
$C:=\{x\in 2^{\omega}\mid G_{x}\text{ is Ramsey positive}\}.$
Proposition 4.1.
The set $C$ is $\mathbf{\Sigma}^{1}_{2}.$
Proof.
Observe that for $x\in 2^{\omega}$ coding $\langle X_{n}\subseteq\mathcal{AR}\mid n<\omega\rangle$, we have $x\in C$ iff
  1. (1)
    for all $n<\omega$, $X_{n}$ is a tree;
  2. (2)
    $\mathcal{R}\setminus\bigcup_{n<\omega}[X_{n}]$ is Ramsey positive.
Note that condition (1) is closed, so it suffices to verify that condition (2) is $\mathbf{\Sigma}^{1}_{2}$. Since $\mathcal{R}\setminus\bigcup_{n<\omega}[X_{n}]$ is analytic, condition (2) holds iff $\mathcal{R}\setminus\bigcup_{n<\omega}[X_{n}]$ contains some non-empty $[s,A]$, i.e.,
$\exists s\in\mathcal{AR}~{}\exists A\in\mathcal{R}\big{(}s\sqsubseteq A\wedge\forall n<\omega~{}[s,A]\cap[X_{n}]=\emptyset\big{)}.$
Since $\mathcal{AR}$ is countable and $\mathcal{R}$ is a Polish space, this condition is $\mathbf{\Sigma}^{1}_{2}$. We conclude that $C$ is $\mathbf{\Sigma}^{1}_{2}$. ∎
5. Conditions for $\mathbf{\Sigma}^{1}_{2}$-completeness
Let $G\subseteq 2^{\omega}\times\mathcal{R}$ and $C\subseteq 2^{\omega}$ be as in Section 4. By Proposition 4.1, $C$ is $\mathbf{\Sigma}^{1}_{2}$. We present sufficient conditions for $C$ to be $\mathbf{\Sigma}^{1}_{2}$-complete.
Given any map $m:\mathcal{AR}\to\omega$, we define $\ell:\mathcal{AR}\to[\omega]^{<\omega}$ and ¯$\overline{\ell}:\mathcal{R}\to[\omega]^{\leq\omega}$ by
$\ell(s):=\{m(t)\mid\emptyset\neq t\sqsubseteq s\}$
and
¯$\overline{\ell}(A):=\{m(t)\mid\emptyset\neq t\sqsubseteq A\}=\bigcup_{n<\omega}\ell(r_{n}(A)).$
Property 5.1.
We consider the following four properties that a map $m:\mathcal{AR}\to\omega$ may satisfy:
  1. (1)
    (Monotonicity) For all $s\sqsubseteq t$, $m(s)\leq m(t)$.
  2. (2)
    (Unboundedness) For all $A$, ¯$\overline{\ell}(A)\in[\omega]^{\omega}$.
  3. (3)
    (Compatibility with $\leq$) There exists $A^{*}$ such that ¯$\overline{\ell}(A^{*})=\omega$ and for all $B\leq A\leq A^{*}$, we have ¯$\overline{\ell}(B)\subseteq\overline{\ell}(A)$.
  4. (4)
    (Selection) For any $s\sqsubseteq A$ and $B^{\prime}\in[\ell(s),\overline{\ell}(A)]$ with |¯$|\overline{\ell}(A)\setminus B^{\prime}|=1$, there is $B\in[s,A]$ such that ¯$\overline{\ell}(B)=B^{\prime}$.
Remark 5.2.
Instead of requiring ¯$\overline{\ell}(A^{*})=\omega$ in the compatibility property, it is sufficient for ¯$\overline{\ell}(A^{*})$ to be maximal in the sense that for every $A\in\mathcal{R}$, we have ¯$\overline{\ell}(A)\subseteq\overline{\ell}(A^{*})$. Given a map $m$ satisfying this modified version of compatibility along with monotonicity, unboundedness, and selection, a slight adjustment will yield a map satisfying all of Property 5.1.
Remark 5.3.
Observe that in the presence of monotonicity, unboundedness, and compatibility, the selection property is equivalent to the following statement:
For any $s\sqsubseteq A$ and $B^{\prime}\in[\ell(s),\overline{\ell}(A)]$, there is $B\in[s,A]$ such that ¯$\overline{\ell}(B)=B^{\prime}$.
Indeed, if ¯$\overline{\ell}(A)\setminus B^{\prime}=\{p_{i}\mid i<\omega\}$ with $p_{0}<p_{1}<\ldots$, we can construct $A\geq B_{0}\geq B_{1}\geq\ldots$ and $s\sqsubseteq s_{0}\sqsubseteq s_{1}\sqsubseteq\ldots$ such that ¯$\overline{\ell}(B_{i})=\overline{\ell}(A)\setminus\{p_{j}\mid j\leq i\}$, $s_{i}\sqsubseteq B_{i}$, and $\ell(s_{i})=\overline{\ell}(B_{i})\cap p_{i+1}$. Then we can see that the limit $B:=\lim_{i}B_{i}\in[s,A]$ exists and ¯$\overline{\ell}(B)=B^{\prime}$.
Note that monotonicity implies $\ell(s)\sqsubseteq\ell(t)$ when $s\sqsubseteq t$, and $\ell(s)\sqsubseteq\overline{\ell}(A)$ when $s\sqsubseteq A$. In addition, Property 5.1 implies that $\ell[\mathcal{AR}]=[\omega]^{<\omega}$ and ¯$\overline{\ell}[\mathcal{R}]=[\omega]^{\omega}$; compatibility and selection are key to this observation.
The following proposition, along with Proposition 4.1, completes the proof of Theorem 1.1.
Proposition 5.4.
Suppose there exists a map $m:\mathcal{AR}\to\omega$ satisfying Property 5.1. Then $C$ is $\mathbf{\Sigma}^{1}_{2}$-hard.
Proof.
Let $H\subseteq 2^{\omega}\times[\omega]^{\omega}$ be a universal $G_{\delta}$ set constructed in the same way as $G$. By [Sab12, Theorem 1], the set $C^{\prime}:=\{x\in 2^{\omega}\mid H_{x}\text{ is Ramsey positive}\}$ is $\mathbf{\Sigma}^{1}_{2}$-complete. Thus, it suffices to construct a continuous reduction from $C^{\prime}$ to $C$.
We define a map $f:\mathcal{P}([\omega]^{<\omega})\to\mathcal{P}(\mathcal{AR})$ as follows. Given $X\subseteq[\omega]^{<\omega}$, define $f(X)\subseteq\mathcal{AR}$ by
$s\in f(X)\iff\ell(s)\in X$
[…]
(*) $A\in[f(X)]\iff\overline{\ell}(A)\in[X]$
for all $A\in\mathcal{R}$.
Define the map $\varphi:2^{\omega}\to 2^{\omega}$ such that if $x\in 2^{\omega}$ codes the sequence $\langle X_{n}\subseteq[\omega]^{<\omega}\mid n<\omega\rangle$, then $\varphi(x)$ codes the sequence $\langle f(X_{n})\subseteq\mathcal{AR}\mid n<\omega\rangle$. Note that $\varphi$ is continuous. We claim that $\varphi$ is a reduction from $C^{\prime}$ to $C$.
First, suppose $x\in C^{\prime}$ and $x$ codes $\langle X_{n}\subseteq[\omega]^{<\omega}\mid n<\omega\rangle$. Since $x\in C^{\prime}$, we must have that each $X_{n}$ is a tree and $[\omega]^{\omega}\setminus\bigcup_{n<\omega}[X_{n}]$ is Ramsey positive. Then each $f(X_{n})$ is a tree, so
$G_{\varphi(x)}=\mathcal{R}\setminus\bigcup_{n<\omega}[f(X_{n})].$
Fix some $s^{\prime}\sqsubseteq A^{\prime}$ such that $[s^{\prime},A^{\prime}]\cap[X_{n}]=\emptyset$ for every $n$. By compatibility and selection, we can find $s\sqsubseteq A\leq A^{*}$ such that $\ell(s)=s^{\prime}$ and ¯$\overline{\ell}(A)=A^{\prime}$. To prove that $G_{\varphi(x)}$ is Ramsey positive, it suffices to show $[s,A]\cap[f(X_{n})]=\emptyset$ for every $n$. Suppose for some $n<\omega$, we can find $B\in[s,A]\cap[f(X_{n})]$. Let $B^{\prime}:=\overline{\ell}(B)$. Note that $B^{\prime}\in[X_{n}]$ by (*5). We have $s^{\prime}\sqsubseteq B^{\prime}$ and $B^{\prime}\subseteq A^{\prime}$ by monotonicity and compatibility, so $B^{\prime}\in[s^{\prime},A^{\prime}]$. But this contradicts that $[s^{\prime},A^{\prime}]\cap[X_{n}]=\emptyset$. So we conclude $[s,A]\cap[f(X_{n})]=\emptyset$ for each $n$, hence $G_{\varphi(x)}$ is Ramsey positive and $\varphi(x)\in C$.
Now suppose we have $x\in 2^{\omega}$ such that $\varphi(x)\in C$ and $x$ codes $\langle X_{n}\subseteq[\omega]^{<\omega}\mid n<\omega\rangle$. Since $\varphi(x)\in C$, every $f(X_{n})$ is a tree and $\mathcal{R}\setminus\bigcup_{n<\omega}[f(X_{n})]$ is Ramsey positive. So we have that every $X_{n}$ is a tree and
$H_{x}=[\omega]^{\omega}\setminus\bigcup_{n<\omega}[X_{n}].$
Fix $s\sqsubseteq A$ such that $[s,A]\cap[f(X_{n})]=\emptyset$ for all $n<\omega$. Let $s^{\prime}:=\ell(s)$ and $A^{\prime}:=\overline{\ell}(A)$. Then $s^{\prime}\sqsubseteq A^{\prime}$ by monotonicity. We claim $[s^{\prime},A^{\prime}]\cap[X_{n}]=\emptyset$ for all $n<\omega$. Suppose for some $n<\omega$, we can find $B^{\prime}\in[s^{\prime},A^{\prime}]\cap[X_{n}]$. By selection, there is $B\in[s,A]$ such that ¯$\overline{\ell}(B)=B^{\prime}$. By (*5), we have $B\in[f(X_{n})]$ since $B^{\prime}\in[X_{n}]$. But this contradicts that $[s,A]\cap[f(X_{n})]=\emptyset$. We conclude that $H_{x}$ is Ramsey positive, thus $x\in C^{\prime}$.
So $\varphi$ is a continuous reduction from $C^{\prime}$ to $C$. Therefore, $C$ is $\mathbf{\Sigma}^{1}_{2}$-hard. ∎
6. Examples
We present a few examples of topological Ramsey spaces where Theorem 1.1 shows that the set of codes for $G_{\delta}$ Ramsey positive sets is $\mathbf{\Sigma}^{1}_{2}$-complete. For each of these spaces, we exhibit a map $m$ witnessing Property 5.1.
6.1. Ellentuck space
The proof of Proposition 5.4 relies on Sabok’s result that the set of codes for $G_{\delta}$ Ramsey positive sets of $[\omega]^{\omega}$ is $\mathbf{\Sigma}^{1}_{2}$-complete, so finding a map $m:[\omega]^{<\omega}\to\omega$ satisfying Property 5.1 is superfluous. We nevertheless provide such a map for the sake of illustration.
Define $m:[\omega]^{<\omega}\to\omega$ by $m(\emptyset):=0$ and $m(s):=\max(s)$ for $s\neq\emptyset$. Then $\ell(s)=s$ and ¯$\overline{\ell}(A)=A$ for all $s\in[\omega]^{<\omega}$ and $A\in[\omega]^{\omega}$. It is clear that monotonicity and unboundedness hold. Letting $A^{*}=\omega$, we see that compatibility holds as well. Finally, selection holds since we can pick $B=B^{\prime}$. So $m$ satisfies Property 5.1.
6.2. Strong subtrees of $2^{<\omega}$
Let $2^{<\omega}$ denote the complete binary tree. We say that $A\subseteq 2^{<\omega}$ is a strong subtree of $2^{<\omega}$ if there exists a set of levels $I\subseteq\omega$ such that
  1. (1)
    $A\subseteq\bigcup_{i\in I}2^{i}$;
  2. (2)
    for every $i\in I$, $A\cap 2^{i}\neq\emptyset$;
  3. (3)
    if $i\in I$, $i^{\prime}=\min\{n\in I\mid n>i\}$, $a\in A\cap 2^{i}$, and $j\in\{0,1\}$, then there is exactly one $b\in A\cap 2^{i^{\prime}}$ extending $a^{\frown}(j)$.
Let $\mathcal{S}_{\infty}(2^{<\omega})$ denote the set of all infinite strong subtrees of $2^{<\omega}$. If $\{i_{k}\mid k<\omega\}$ is the increasing enumeration of the levels of $A$, define $r_{n}(A)=A\cap\bigcup_{k<n}2^{i_{k}}$ for each $n<\omega$. Then $(\mathcal{S}_{\infty}(2^{<\omega}),\subseteq,r)$ is a topological Ramsey space; see [Mil81] and [Tod10, Section 6.1] for details.
Let $\mathcal{S}_{<\infty}(2^{<\omega})=\mathcal{A}\mathcal{S}_{\infty}(2^{<\omega})$ denote the set of finite strong subtrees of $2^{<\omega}$. Define $m:\mathcal{S}_{<\infty}(2^{<\omega})\to\omega$ by
$m(s):=\begin{cases}0,&\quad\text{if }s=\emptyset\\ \max\{n\mid s\cap 2^{n}\neq\emptyset\},&\quad\text{otherwise.}\end{cases}$
Observe that for any $s\in\mathcal{S}_{<\infty}(2^{<\omega})$ and $A\in\mathcal{S}_{\infty}(2^{<\omega})$, $\ell(s)=\{n\mid s\cap 2^{n}\neq\emptyset\}$ and ¯$\overline{\ell}(A)=\{n\mid A\cap 2^{n}\neq\emptyset\}$ are the corresponding sets of levels. Monotonicity and unboundedness are clear, and compatibility holds with $A^{*}=2^{<\omega}$. Selection also holds: intuitively, finding an appropriate $B\in[s,A]$ given $B^{\prime}=\overline{\ell}(A)\setminus\{k\}\in[\ell(s),\overline{\ell}(A)]$ amounts to thinning the nodes in $A$ at levels above $k$.
Thus, $m$ satisfies Property 5.1, and Theorem 1.1 implies that the set of codes for $G_{\delta}$ Ramsey positive subsets of $\mathcal{S}_{\infty}(2^{<\omega})$ is $\mathbf{\Sigma}^{1}_{2}$-complete.
6.3. Infinite sequences of words with variables
Let $L=\bigcup_{n<\omega}L_{n}$ be a set, where each $L_{n}$ is finite and $L_{n}\subseteq L_{n+1}$. Let $v\notin L$. Define $W_{Lv}$ to be the set of all words—non-empty finite strings of elements from $L\cup\{v\}$—in which $v$ appears. Define
$W_{Lv}^{[\infty]}:=\{\langle a_{n}\rangle_{n<\omega}\mid\forall n<\omega~{}\big{(}a_{n}\in W_{Lv}\wedge|a_{n}|>\sum_{i<n}|a_{i}|\big{)}\}.$
Given $A\in W_{Lv}^{[\infty]}$, let
$[A]_{Lv}:=\{a_{n_{0}}[\lambda_{0}]^{\frown}\ldots^{\frown}a_{n_{k}}[\lambda_{k}]\in W_{Lv}\mid n_{0}<\ldots<n_{k}\wedge\forall 0\leq i\leq k~{}\lambda_{i}\in L_{n_{i}}\cup\{v\}\},$
where $a[\lambda]$ denotes the string obtained by replacing every instance of $v$ in $a$ with $\lambda$. If $a=a_{n_{0}}[\lambda_{0}]^{\frown}\ldots^{\frown}a_{n_{k}}[\lambda_{k}]\in[A]_{Lv}$, define
$\text{supp}_{A}(a):=\{n_{0},\ldots,n_{k}\}.$
Note that the condition on the lengths of the words in $A$ guarantees that $\text{supp}_{A}(a)$ is well-defined.
For $A=\langle a_{n}\rangle_{n<\omega},B=\langle b_{n}\rangle_{n<\omega}\in W_{Lv}^{[\infty]}$, define $A\leq B$ iff
  1. (i)
    for all $n<\omega$, $a_{n}\in[B]_{Lv}$;
  2. (ii)
    for all $k<n$, $\max(\text{supp}_{B}(a_{k}))<\min(\text{supp}_{B}(a_{n}))$.
For $k<\omega$ and $A=\langle a_{n}\rangle_{n<\omega}\in W_{Lv}^{[\infty]}$, define $r_{k}(A):=\langle a_{n}\rangle_{n<k}$. Then $(W_{Lv}^{[\infty]},\leq,r)$ is a topological Ramsey space; see [Car88] and [Tod10, Section 5.3] for details.
Let $W_{Lv}^{[<\infty]}:=\mathcal{A}W_{Lv}^{[\infty]}$. Define $m:W_{Lv}^{[<\infty]}\to\omega$ by $m(\emptyset):=0$ and $m(\langle a_{i}\rangle_{i<n}):=\lfloor\log_{2}|a_{n-1}|\rfloor$ if $n\geq 1$. It is clear that monotonicity and unboundedness are satisfied. We can see that compatibility holds if we let $A^{*}=\langle a_{n}^{*}\rangle_{n<\omega}$, where $a_{n}^{*}$ is the word $vv\ldots v$ consisting of $v$ exactly $2^{n}$ times. We remark that although we have ¯$\overline{\ell}(B)\subseteq\overline{\ell}(A)$ when $B\leq A\leq A^{*}$, this is not true for all $B\leq A$. Finally, to show selection, suppose $s\sqsubseteq A$ and $B^{\prime}=[\ell(s),\overline{\ell}(A)]$ with ¯$\overline{\ell}(A)\setminus B^{\prime}=\{k\}$. If $A=\langle a_{n}\rangle_{n<\omega}$, we can form $B\in[s,A]$ with ¯$\overline{\ell}(B)=B^{\prime}$ by setting $B=\langle a_{n}\mid n<\omega,\lfloor\log_{2}|a_{n}|\rfloor\neq k\rangle$.
We conclude that $m$ satisfies Property 5.1, hence the set of codes for $G_{\delta}$ Ramsey positive subsets of $W_{Lv}^{[<\infty]}$ is $\mathbf{\Sigma}^{1}_{2}$-complete by Theorem 1.1.
6.4. High-dimensional Ellentuck spaces
For $k\geq 2$, we define the topological Ramsey space $\mathcal{E}_{k}$ as in [Dob16a]. Let $\omega^{\not\downarrow\leq k}$ denote the set of all non-decreasing sequences of natural numbers of length at most $k$, i.e.,
$\omega^{\not\downarrow\leq k}=\{\langle u_{0},u_{1},\ldots,u_{p-1}\rangle\mid 0\leq p\leq k\wedge u_{0}\leq u_{1}\leq\ldots\leq u_{p-1}\}.$
Let $\omega^{\not\downarrow k}$ denote the set of all non-decreasing sequences of natural numbers of length exactly $k$.
Define the well-order $\prec$ on $\omega^{\not\downarrow\leq k}$ such that
  1. (1)
[…]
  2. (2)
    for $\langle u_{0},\ldots,u_{p-1}\rangle,\langle v_{0},\ldots,v_{q-1}\rangle\in\omega^{\not\downarrow\leq k}$ with $p,q>0$, $\langle u_{0},\ldots,u_{p-1}\rangle\prec\langle v_{0},\ldots,v_{q-1}\rangle$ iff either
    1. (a)
      $u_{p-1}<v_{q-1}$, or
    2. (b)
      $u_{p-1}=v_{q-1}$ and $\langle u_{0},\ldots,u_{p-1}\rangle<_{\text{lex}}\langle v_{0},\ldots,v_{q-1}\rangle$.
Then with respect to $\prec$, $\omega^{\not\downarrow\leq k}$ has order type $\omega$. For $a<\omega$, let $\vec{j}_{a}$ denote the $a^{\text{th}}$ element of $(\omega^{\not\downarrow\leq k},\prec)$. Given any $\vec{u}\in\omega^{\not\downarrow\leq k}$, define $a_{\vec{u}}$ to be the unique $a<\omega$ such that $\vec{u}=\vec{j}_{a}$. For $b<\omega$, let $\vec{i}_{b}$ denote the $b^{\text{th}}$ element of $\omega^{\not\downarrow k}$ with respect to the order $\prec$, inherited from $\omega^{\not\downarrow\leq k}$.
We define a function $\widehat{W_{k}}:\omega^{\not\downarrow\leq k}\to[\omega]^{\leq k}$ as follows. For $\vec{u}\in\omega^{\not\downarrow\leq k}$ with $|\vec{u}|=p$, define $\widehat{W_{k}}(\vec{u}):=\{a_{\vec{u}\mathord{\upharpoonright}q}\mid 1\leq q\leq p\}$. Note that
$a_{\vec{u}\mathord{\upharpoonright}1}<a_{\vec{u}\mathord{\upharpoonright}2}<\ldots<a_{\vec{u}\mathord{\upharpoonright}p}=a_{\vec{u}},$
so $\widehat{W_{k}}(\vec{u})\in[\omega]^{p}$. Let $W_{k}:=\widehat{W_{k}}\mathord{\upharpoonright}\omega^{\not\downarrow k}$. Define
$\mathbb{W}_{k}:=W_{k}[\omega^{\not\downarrow k}]\subseteq[\omega]^{k}$
and
$\widehat{\mathbb{W}_{k}}:=\widehat{W_{k}}[\omega^{\not\downarrow\leq k}]\subseteq[\omega]^{\leq k}.$
Observe that $\widehat{\mathbb{W}_{k}}$ is the tree formed from all initial segments of elements of $\mathbb{W}_{k}$.
A function $\widehat{A}:\omega^{\not\downarrow\leq k}\to\widehat{\mathbb{W}_{k}}$ is an $\mathcal{E}_{k}$-tree if it satisfies the following conditions:
  1. (i)
    for all $a<\omega$, $\widehat{A}(\vec{j}_{a})\in[\omega]^{|\vec{j}_{a}|}$;
  2. (ii)
    for $1\leq a<\omega$, $\max(\widehat{A}(\vec{j}_{a}))<\max(\widehat{A}(\vec{j}_{a+1}))$;
  3. (iii)
    for $a,b<\omega$, $\widehat{A}(\vec{j}_{a})\sqsubseteq\widehat{A}(\vec{j}_{b})$ iff $\vec{j}_{a}\sqsubseteq\vec{j}_{b}$.
Given an $\mathcal{E}_{k}$-tree $\widehat{A}$, define
$[\widehat{A}]:=\widehat{A}\cap(\omega^{\not\downarrow k}\times\mathbb{W}_{k})=\widehat{A}\mathord{\upharpoonright}\omega^{\not\downarrow k}.$
Define
[…]
Note that condition (ii) guarantees that every $A\in\mathcal{E}_{k}$ is uniquely determined by $A[\omega^{\not\downarrow k}]=\{A(\vec{i}_{n})\mid n<\omega\}$. Furthermore, there is a unique $\mathcal{E}_{k}$-tree $\widehat{A}$ for which $[\widehat{A}]=A$ by conditions (i) and (iii). For $A,B\in\mathcal{E}_{k}$, we define $A\leq B$ iff $A[\omega^{\not\downarrow k}]\subseteq B[\omega^{\not\downarrow k}].$ For $n<\omega$ and $A\in\mathcal{E}_{k}$, let $r_{n}(A):=A\mathord{\upharpoonright}\{\vec{i}_{p}\mid p<n\}$. By [Dob16a, Theorem 3.17], $(\mathcal{E}_{k},\leq,r)$ is a topological Ramsey space.
Where convenient, we will write $\widehat{A}(u_{0},u_{1},\ldots,u_{p-1})$ instead of $\widehat{A}(\langle u_{0},u_{1},\ldots,u_{p-1}\rangle)$ below. Define $m:\mathcal{A}\mathcal{E}_{k}\to\omega$ as follows. Let $m(\emptyset):=0$. Given $s$ with domain $\{\vec{i}_{n}\mid n<p\}$ for some $p>0$, let $M_{s}:=\max\{M\mid\exists n<p~{}\langle M\rangle\sqsubseteq\vec{i}_{n}\}$. Note that there must be $n_{s}<p$ such that $\vec{i}_{n_{s}}=\langle M_{s},M_{s},\ldots,M_{s}\rangle$. Define $m(s):=\min(s(\vec{i}_{n_{s}}))$.
We check that $m$ satisfies Property 5.1. To see monotonicity, note that if $\emptyset\neq s\sqsubseteq t$, then we have $M_{s}\leq M_{t}$, $\langle M_{s}\rangle\preceq\langle M_{t}\rangle$, and $s(\vec{i}_{n_{s}})=t(\vec{i}_{n_{s}})$. Fix any $A\in\mathcal{E}_{k}$ with $t\sqsubseteq A$. Then
$m(s)=\min(t(\vec{i}_{n_{s}}))=\max(\widehat{A}(M_{s}))\leq\max(\widehat{A}(M_{t}))=\min(t(\vec{i}_{n_{t}}))=m(t),$
so monotonicity holds.
It is not difficult to see that ¯$\overline{\ell}(A)=\{\max(\widehat{A}(n))\mid n<\omega\}$ for all $A\in\mathcal{E}_{k}$. Since $\max(\widehat{A}(0))<\max(\widehat{A}(1))<\ldots$ by condition (ii), we have ¯$\overline{\ell}(A)\in[\omega]^{\omega}$, hence unboundedness is satisfied.
Observe that $m$ does not satisfy compatibility as written. We will instead show that $m$ satisfies the modified version of compatibility given in Remark 5.2. If we set $A^{*}=W_{k}$, then $m(A^{*})$ is maximal. It suffices to verify that $B\leq A\leq A^{*}$ implies ¯$\overline{\ell}(B)\subseteq\overline{\ell}(A)$. Note that
¯$\overline{\ell}(B)=\{\max(\widehat{B}(n))\mid n<\omega\}\subseteq\{\max(\widehat{A}(n))\mid n<\omega\}=\overline{\ell}(A)$
whenever $B\leq A$, so $m$ satisfies the modified version of compatibility.
Finally, we show that $m$ satisfies selection. Consider $s\sqsubseteq A$ and $B^{\prime}\in[\ell(s),\overline{\ell}(A)]$ with ¯$\overline{\ell}(A)\setminus B^{\prime}=\{x\}$. Fix $u<\omega$ such that $\widehat{A}(u)=\{x\}$. Define $\widehat{B}:\omega^{\not\downarrow\leq k}\to\widehat{\mathbb{W}_{k}}$ as follows. Consider any $\langle u_{0},u_{1},\ldots,u_{p-1}\rangle\in\omega^{\not\downarrow\leq k}$. For $q<p$, define
$v_{q}:=\begin{cases}u_{q},&\quad\text{if }u_{q}\leq u\text{ and }u_{0}\neq u\\ u_{q}+1,&\quad\text{otherwise}.\end{cases}$
Let $\widehat{B}(u_{0},u_{1},\ldots,u_{p-1}):=\widehat{A}(v_{0},v_{1},\ldots,v_{p-1})$. We can verify that $\widehat{B}$ is an $\mathcal{E}_{k}$-tree, so we have $B:=[\widehat{B}]\in\mathcal{E}_{k}$. By definition of $B$, we have $B\leq A$. Note that since $x>\max(\ell(s))$ and $B(\vec{i})=A(\vec{i})$ for all $\vec{i}\in\omega^{\not\downarrow k}$ with $\vec{i}\prec\langle u\rangle$, we must have $s\sqsubseteq B$. So $B\in[s,A]$. Since $\widehat{B}(v)=\widehat{A}(v)$ for all $v<u$ and $\widehat{B}(v)=\widehat{A}(v+1)$ for $v\geq u$, we have
¯$\overline{\ell}(B)=\overline{\ell}(A)\setminus\{\max(\widehat{A}(u))\}=\overline{\ell}(A)\setminus\{x\}=B^{\prime}.$
Thus, $m$ satisfies Property 5.1 with the modification of compatibility according to Remark 5.2, hence Theorem 1.1 applies.
6.5. Other examples
We note that we can find a map satisfying Property 5.1 for the following spaces: $\text{FIN}_{k}^{[\infty]}$ as defined in [Tod10, Section 5.2]; $\mathcal{E}_{\infty}$ as defined in [Tod10, Section 5.6]; $\text{FIN}_{Lv}^{[\infty]}$ as defined in [Tod10, Section 5.3]; $\mathcal{M}_{\infty}$ as defined in [Tod10, Section 5.7]; $\mathcal{R}_{1}$ as defined in [DT14]. So Theorem 1.1 applies to all of these spaces.
There are several generalizations of spaces mentioned above where we have not been able to verify the existence of a map satisfying Property 5.1. In [Dob16b], Dobrinen defined spaces $\mathcal{E}_{B}$ for uniform barriers on $\omega$ as an extension of the spaces $\mathcal{E}_{k}$ from [Dob16a]. We suspect that a map similar to the $m$ defined in Section 6.4 may satisfy Property 5.1 in $\mathcal{E}_{B}$; however, we have not been able to prove it. Likewise, we do not know if we can apply Theorem 1.1 to the spaces $\mathcal{R}_{\alpha}$, $\alpha<\omega_{1}$, defined in [DT15] as an extension of the space $\mathcal{R}_{1}$ defined in [DT14].
7. Further questions
For all the topological Ramsey spaces we have considered so far, we either have a map satisfying Property 5.1 or have a plausible candidate map. We ask the following related questions.
Problem 1.
Do all topological Ramsey spaces $\mathcal{R}$ admit a map $\mathcal{AR}\to\omega$ satisfying Property 5.1?
Problem 2.
Is there a topological Ramsey space whose set of codes for $G_{\delta}$ Ramsey positive sets is not $\mathbf{\Sigma}^{1}_{2}$-complete?
In particular, we are interested to know whether a map $m:\mathcal{AR}\to\omega$ satisfying Property 5.1 can be produced solely using the assumptions A.1-A.4 from [Tod10, Section 5.1].